# Teleportation=Translation: Continuous recovery of black hole information

### v3

@@ Acknowledgements.
I acknowledge the assistance of Large Language Models, specifically Gemini and Chatgpt, in drafting the text, providing valuable feedback, and identifying errors. I take full intellectual responsibility for the content of this work. This work is supported by the National Research Foundation of Korea(NRF) grant with grant number RS-2023-00208047, NRF-2020R1A6A1A03047877, RS-2025-00553127. APPENDIX:
@@ Appendix A Notation table
$\tilde{\mathcal{M}},\ \tilde{\mathcal{N}}$ semifinite crossed-product algebras (envelopes)
@@
$\tilde{E}$, $\tilde{E}_{s}$ lifted conditional expectation and its idempotent interpolants
@@
$U_{\Gamma}$, $U_{\tilde{\Gamma}}$ discrete and lifted canonical shift unitaries
$\quad\tilde{\mathcal{A}}$ relative commutant $\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ (island algebra)
$\tilde{\mathcal{M}}_{1}(s)$ basic extension algebra $\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$
$\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$ accessible radiation sector (basic extension’s commutant)
$\tilde{U}(s)$, $\tilde{G}$ continuous unitary path and its self-adjoint generator
$h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ non-commutative Radon–Nikodym derivative
$K_{\mathcal{X}}$, $P$ modular Hamiltonian of $\mathcal{X}$ and modular momentum (difference of Hamiltonians)
@@ Appendix B Finite-Dimensional Model: Idempotency Failure and Unitary Interpolation
At the midpoint $s=0.5$, the map becomes $E_{0.5}(A)=0.5A+0.25\mathrm{Tr}(A)\mathbf{1}$. Applying the map again to test idempotency yields:
@@
Since $E_{0.5}^{2}(A)\neq E_{0.5}(A)$, this demonstrates the fatal failure of static idempotency for the naive path. This failure is precisely why the Haagerup–Kosaki lift is required in the Type III setting.
@@
Thus, the endpoint coincides with the swap up to a global phase factor of i-i. This illustrates the physical goal of the construction: finding a continuous evolution $U(s)$ that connects the initial state to the final shift. This unitary path serves as a consistent alternative to the naive CP-map interpolation, which was shown to be mathematically problematic in Example B.1.
@@ Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of Type III von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\widehat{\mathcal{N}}^{+}$.
Suppose we select a faithful normal weight $\omega$ on $\mathcal{M}$ that is compatible with the inclusion, meaning its modular automorphism group $\sigma_{t}^{\omega}$ leaves the subalgebra invariant:
@@
(This condition implies that $\omega$ can be written as $\omega=\psi\circ\mathcal{E}$ for some faithful normal weight $\psi$ on $\mathcal{N}$).
Under this modular covariance assumption, the crossed product construction yields the following structure:
@@
    $\tau\circ\theta_{t}=e^{-t}\tau.$ (42)
    Furthermore, the connection to the base algebra is maintained by the canonical OVW $\underline{\mathcal{E}}:\tilde{\mathcal{M}}^{+}\to\widehat$$\mathcal{M}$$^{+}$ (often denoted as the dual weight construction), formally given by:
@@
    $\theta_{t}(\tilde{\mathcal{N}})=\tilde{\mathcal{N}}$. (44)
@@
    which is trace-preserving (i.e., $\tau\circ\tilde{E}=\tau$) and is equivariant with respect to the dual action $\theta_{t}$ (i.e., $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$).
Proof.
The proof formalizes the conversion of the Type III unbounded structure into a Type $\text{II}_{\infty}$ bounded projection.
@@
    Compatibility: The assumption $\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}$ ensures that the crossed product of the subalgebra, $\tilde{\mathcal{N}}$, is naturally a subalgebra of $\tilde{\mathcal{M}}$. Without this modular covariance, the relationship between the lifted algebras is not well-defined.
@@
    Semifinite Structure and Trace: The theory of crossed products guarantees that $\tilde{\mathcal{M}}$ is a semifinite algebra equipped with a canonical trace $\tau$. The scaling property (42) is a defining characteristic of the Type II crossed product derived from a Type III algebra, allowing us to treat modular dynamics as trace-scaling automorphisms. Consequently, the existence of this trace allows for the construction of conditional expectations that are not available in the Type III setting. ă
@@
    Construction of $\tilde{E}$: The existence of the trace-preserving conditional expectation $\tilde{E}$ follows from the compatibility of the dual weights on $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Specifically, because the original weights were chosen to be compatible ($\omega=\psi\circ\mathcal{E}$), their dual traces on the crossed products coincide, implying the existence of a conditional expectation. This $\tilde{E}$ effectively renormalizes the original OVW $\mathcal{E}$ into a norm-one projection on the semifinite algebra. ă
@@
    Equivariance: The construction explicitly preserves the structure of the dual action $\theta_{t}$, ensuring $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$. This property is crucial for proving that the analytic domains are preserved under the operation of $\tilde{E}$.

### v4

@@ Acknowledgements.
I acknowledge the assistance of Large Language Models, specifically Gemini and Chatgpt, in drafting the text, providing valuable feedback, and identifying errors. I take full intellectual responsibility for the content of this work. This work is supported by the National Research Foundation of Korea(NRF) grant with grant number RS-2019-NR040081, RS-2023-00208047, RS-2025-00553127. APPENDIX:
@@ Appendix A Notation table
$\tilde{\mathcal{M}},\ \tilde{\mathcal{N}}$ semifinite crossed-product algebras (lifted, enveloped algebras)
$\quad\mathcal{X}^{\prime}$ commutant of algebra $\mathcal{X}$
@@
$\quad s,\ p$ interpolated and $L^{p}$-space parameters
$L^{p}(\mathcal{X})$ non-commutative $L^{p}$ spaces associated with $\mathcal{X}$
@@
$\mathcal{E},\ T$ operator-valued weight (OVW) and dual OVW
$\tilde{\mathcal{M}}_{1}(s),\ \tilde{\mathcal{N}}(s)$ basic extension algebra $\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$, interpolated subalgebra
@@
$U_{\Gamma},\ \tilde{U}_{\Gamma},\ \tilde{U}(s)$ discrete and lifted canonical shift unitaries
$D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$ Connes-Takesaki Radon-Nikodym cocycle
$h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ Radon–Nikodym derivative
$\quad\tilde{G}$ self-adjoint generator
$K_{\mathcal{X}},\ P$ modular Hamiltonian and modular momentum
@@ Appendix B Finite-Dimensional Model: Idempotency Failure and Unitary Interpolation
At the midpoint $s=0.5$, the map becomes $E_{0.5}(A)=0.5A+0.25\mathrm{Tr}(A)\mathbf{1}$. Applying the map again to test idempotency, and noting that $E_{0.5}(\mathbf{1})=\mathbf{1}$, we obtain:
@@
Since $E_{0.5}^{2}(A)\neq E_{0.5}(A)$, this demonstrates the breakdown of the idempotency property for the naive path. This failure illustrates why the more sophisticated Haagerup–Kosaki lift is required to preserve algebraic structure in the Type III setting.
@@
Thus, the endpoint coincides with the swap operation up to a global phase factor of i-i. This illustrates the physical goal of our construction: finding a continuous evolution $U(s)$ that connects the initial state to the final shift. While this finite-dimensional model implies trivial modular dynamics (and thus does not exhibit $\tilde{G}=2P$), it effectively demonstrates the necessity of a unitary path as a consistent alternative to the naive CP-map interpolation shown in Example B.1.
@@ Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of Type III von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\widehat{\mathcal{N}}^{+}$. Let $\omega$ be a faithful normal weight on $\mathcal{M}$ that is compatible with the inclusion, meaning its modular automorphism group $\sigma_{t}^{\omega}$ leaves the subalgebra invariant:
@@
This condition implies that $\omega$ decomposes as $\omega=\psi\circ\mathcal{E}$ for some faithful normal weight $\psi$ on $\mathcal{N}$. Under this modular covariance assumption, the crossed product construction yields the following structure:
@@
    $\tau\circ\theta_{q}=e^{-q}\tau.$ (48)
    Furthermore, the connection to the base algebra is maintained by the canonical dual OVW $T:\tilde{\mathcal{M}}^{+}\to\widehat{\mathcal{M}}^{+}$, formally given by the integral of the dual action:
@@
    $\theta_{q}(\tilde{\mathcal{N}})=\tilde{\mathcal{N}}$. (50)
@@
    which is trace-preserving (i.e., $\tau\circ\tilde{E}=\tau$) and is equivariant with respect to the dual action $\theta_{q}$ (i.e., $\tilde{E}\circ\theta_{q}=\theta_{q}\circ\tilde{E}$).
Outline of Construction.
The construction resolves the Type III pathology by embedding the algebra into a larger Type $\text{II}_{\infty}$ system. The key steps are:
@@
    Compatibility: The assumption $\sigma_{t}^{\omega}(\mathcal{N})\subset\mathcal{N}$ ensures that the modular dynamics of the ambient algebra do not mix the subalgebra with the complement, allowing $\tilde{\mathcal{N}}$ to be a well-defined subalgebra of $\tilde{\mathcal{M}}$.
@@
    Semifinite Structure: The theory of crossed products implies that $\tilde{\mathcal{M}}$ is semifinite. The scaling property (48) is the defining characteristic of a Type II crossed product derived from a Type III algebra, allowing us to treat modular automorphisms as trace-scaling operations.
@@
    Construction of $\tilde{E}$: The existence of the trace-preserving conditional expectation $\tilde{E}$ follows from the compatibility of the dual weights. Since the original weights satisfy $\omega=\psi\circ\mathcal{E}$, their associated dual traces on the crossed products are consistent. This consistency allows OVW $\mathcal{E}$ (which was unbounded) to be renormalized into a contractive projection $\tilde{E}$ on the semifinite algebra.
@@
    Equivariance: By construction, $\tilde{E}$ respects the dual action $\theta_{q}$. This equivariance is crucial for defining the analytic domains required for the interpolation $E_{s}$, ensuring that the path remains within the correct operator spaces.